# Kai Shu, Lu Cheng, Huan Liu

### v4

@@ 5.3.2 Retrieval-Augmented Generation (RAG)
Recent developments in Retrieval-Augmented Generation (RAG) have explored the capacity of LLMs to self-evaluate and self-improve without the need for annotated datasets or parameter adjustments Chen et al. (2024d). Li and Qiu (2023) introduce the Memory-of-Thought (MoT) framework, a two-stage self-reflective model that enhances an LLM’s reasoning abilities autonomously. In the first stage, the model generates high-confidence reasoning on an unlabeled dataset, storing this as memory. During the testing phase, the model recalls this memory by judging the relevance of each of them with the current question and choose the most relevant ones as demonstration.
@@ 5.4.1 Reasoning Path Selection
Additionally, there are also many works that focus on step-level reasoning path selection, leveraging the judge LLM as the process reward model (PRM) to evaluate state scores. Creswell et al. (2023) decompose the reasoning process into Selection and Inference. In the selection step, they leverage the LLM itself to judge and evaluate each potential reasoning trace, selecting the appropriate one for the following inference step. Xie et al. (2024b) propose the Kwai-STaR framework, which transforms LLMs into state-transition reasoners to judge and select the best reasoning state for themselves in mathematical reasoning. Lightman et al. (2023) train a LLM as the PRM to conduct inference-time supervision and perform best-of-N sampling strategy during reasoning stage. Following them, Setlur et al. (2024) further propose process advantage verifiers (PAVs), generating rewards based on the changes in the likelihood of producing correct responses in the future. Other works simulate advanced cognitive structures as the reasoning process. Hao et al. (2023) employ LLMs as a world model to simulate the state of the environment and perform Monte Carlo Tree Search (MCTS) to improve performance on tasks requiring deliberate path selection. Furthermore, Besta et al. (2024) takes the outputs generated by LLMs as an arbitrary graph. An LLM thought is modeled as a vertex, while an edge is a dependency between thoughts. This framework enables a systematic judge of coherence and logical reasoning for each reasoning state. Yao et al. (2023a) propose Tree-of-Thoughts (ToT), where each thought serves as an intermediate step toward problem-solving. It decomposes reasoning into steps, self-evaluates and judges progress at each state, and uses search algorithms with LMs to judge thought paths through lookahead and backtracking. Besides, there are also studies train critique-based LLM judges Ankner et al. (2024); Yu et al. (2024b); Wang et al. (2024c); Lan et al. ; Xie et al. (2024b); Yao et al. (2024) which provide fine-grained verbal feedback to boost the reasoning process.
@@
DHP Wang et al. (2024g) Various Tasks 400 Human Discernment Score General Performance
EvalBiasBench Park et al. (2024) Alignment 80 Human Accuracy Various Bias
@@
R-Judge Yuan et al. (2024b) Multi-turn Interaction 569 Human F1, Recall, Spec, Effect Safety
@@ 6.1 General Performance
  • JUDGE-BENCH Tan et al. (2024a), DHP Wang et al. (2024g), RewardBench Lambert et al. (2024), SOS-BENCH Penfever et al. (2024) and JuStRank Gera et al. (2024), which operate at larger scales, utilizing metrics like Cohen’s kappa, Discernment Score, and normalized accuracy to benchmark general LLM performance.
@@ 6.3 Challenging Task Performance
Benchmarks designed for difficult tasks push the boundaries of LLM evaluation. For example, Arena-Hard Auto Li et al. (2024k) and JudgeBench Tan et al. (2024a) select harder questions based on LLMs’ performance for conversational QA and various reasoning tasks, respectively. CALM Ye et al. (2024a) explores alignment and challenging scenarios, using metrics such as separability, agreement, and hacked accuracy to evaluate performance in manually identified hard data sets.
@@ 6.4 Domain-Specific Performance
Domain-specific benchmarks provide task-focused evaluations to assess LLMs’ effectiveness in specialized contexts. Concretely, Raju et al. (2024) measure separability and agreement across tasks, leveraging metrics such as Brier scores for insights in specific domains such as coding, medical, finance, law and mathematics. CodeJudge-Eval Zhao et al. (2024a) specifically evaluates LLMs for judging code generation with execution-focused metrics such as accuracy and F1 score. This idea has also been adopted by several following works in code summarization and generation evaluation Wu et al. (2024b); Yang et al. (2024); Tong and Zhang (2024); Lin et al. (2024); Gomes et al. (2024); Yu et al. (2024d)
@@ 7.1 Bias & Vulnerability
In addition to these general biases, specific evaluative biases emerge when LLMs act as judges. Order Bias is a prominent issue where the sequence of candidates influences preferences Zheng et al. (2023); Wang et al. (2023c); Koo et al. (2023a); Shi et al. (2024b). This bias can distort evaluation outcomes, particularly in pairwise comparisons, and is more pronounced when the quality gap between competing responses is small Shi et al. (2024b); Wang et al. (2023c). Egocentric Bias arises when LLMs favor outputs generated by the same model, compromising objectivity Liu et al. (2023c); Koo et al. (2023a); Wataoka et al. (2024); Panickssery et al. (2024). This issue is particularly pronounced when evaluation metrics are designed using the same model, leading to inflated scores for outputs from the originating model Liu et al. (2023c). Length Bias, another prevalent challenge, skews evaluations by disproportionately favoring longer or shorter responses regardless of quality Zheng et al. (2023); Koo et al. (2023a). Additional biases, such as Misinformation Oversight Bias, Authority Bias, and Beauty Bias, further complicate LLM evaluations. For instance, misinformation oversight bias reflects a tendency to overlook factual errors, authority bias favors statements from perceived authoritative sources, and beauty bias prioritizes visually appealing content over substantive quality Chen et al. (2024b, d); Stephan et al. (2024). Similarly, Verbosity Bias shows a preference for lengthier explanations, often equating verbosity with quality, which can mislead the judgment process Yuan et al. (2024a). Furthermore, Sentiment Bias skews evaluations based on emotional tone, favoring responses with positive phrasing Ye et al. (2024a).
@@
Future Direction. One promising direction for future research is to integrate Retrieval-Augmented Generation (RAG) frameworks into LLM evaluation processes Chen et al. (2024d). By combining generative and retrieval capabilities, these frameworks can reduce biases such as self-preference and factuality issues by grounding evaluations in external, verifiable data sources. Another promising avenue is the use of bias-aware datasets, like OFFSETBIAS, to systematically address inherent biases in LLM-as-a-judge systems Park et al. (2024). Incorporating such datasets into training pipelines allows LLMs to better distinguish superficial qualities from substantive correctness, thereby enhancing fairness and reliability. Exploring fine-tuned LLMs as scalable judges, exemplified by the JudgeLM framework, represents another intriguing direction Zhu et al. (2023). Techniques like swap augmentation and reference support embedded in this framework can systematically mitigate biases, improve evaluation consistency, and extend the applicability of LLM-based judgments to open-ended tasks. Additionally, advancing zero-shot comparative assessment frameworks offers significant promise Liusie et al. (2023). These frameworks can refine pairwise comparison techniques and implement debiasing strategies, improving fairness and reliability across diverse evaluation domains without the need for extensive prompt engineering or fine-tuning. Finally, methods like JudgeDeceiver-resistant calibration and adversarial phrase detection strategies need further exploration to secure LLM-as-a-judge frameworks from attacks Shi et al. (2024a); Raina et al. (2024); Zhou et al. (2024a); Li et al. (2024d).
@@ 7.2 Dynamic & Complex Judgment
Earlier works about LLM-as-a-judge usually adopt static and straightforward approaches that directly prompt the judge LLM to perform the assessment Zheng et al. (2023). Recently, more dynamic and complex judgment pipelines have been proposed to address various limitations, improving the robustness and effectiveness of LLM-as-a-judge. One approach in direction is to follow the concept of “LLM-as-a-examiner”, where the system dynamically and interactively generates both questions and judgments based on the candidate LLMs’ performance Yu et al. (2024c); Bai et al. (2023a); Kim et al. (2024a). Other works focus on making judgments based on two or more candidate LLMs battle and debating results Moniri et al. (2024); Zhao et al. (2024c); Kenton et al. . These dynamic judgment methods largely improve the judge LLM’s understanding of each candidate and potentially prevent the data contamination problem Deng et al. (2024); Golchin and Surdeanu in LLMs’ evaluation. Additionally, building complex and complicated judgment pipelines or agents is another popular research area Li et al. (2023a); Chan et al. (2023); Zhuge et al. (2024); Zheng et al. (2024). These methods typically involve multi-agent collaboration, along with well-designed planning and memory systems, enabling the judge LLMs to handle more complex and diverse judgment scenarios.
Future Direction. One promising direction for future research is to equip LLMs with human-like judgment capabilities Yuan et al. (2024a); Liang et al. (2024b). These designs can borrow insights from human behavior when making a judgment, such as anchoring and comparing, hindsight and reflection, and meta-judgment. Another intriguing avenue would be to develop an adaptive difficulty assessment system using LLMs Hu (2024). This system would adjust the difficulty of questions based on the current performance of the candidates. Such an adaptive and dynamic system could address a significant limitation in LLM evaluation, as static benchmarks often fail to accurately assess LLMs with varying capabilities.
@@ 7.3 Self-Judging
LLM-based evaluators, such as GPT-4, are widely used for assessing outputs but face significant challenges, particularly Egocentric Bias, where models favor their own responses over those from external systems Liu et al. (2023b); Bai et al. (2023a); Zheng et al. (2023). This self-preference undermines impartiality and creates a “chicken or the egg” dilemma: robust evaluators are essential for developing powerful LLMs, yet advancing LLMs depends on unbiased evaluators. Other issues include Self-Enhancement Bias, where models overrate their own outputs Li et al. (2023a), and Reward Hacking, where over-optimization of specific signals leads to less generalizable evaluations Wu et al. (2024a). Additionally, reliance on Static Reward Models limits adaptability, while biases like Positional and Verbosity Biases distort judgments by favoring response order or length over quality Yuan et al. (2024c); Wang et al. (2024f). The high cost and limited scalability of human annotations further complicate the creation of dynamic and reliable evaluation systems Srivastava et al. (2022); Liang et al. (2022).
@@
Future Direction. One promising direction for future research is the development of collaborative evaluation frameworks like Peer Rank and Discussion (PRD) Li et al. (2023a). These frameworks leverage multiple LLMs to collectively evaluate outputs, using weighted pairwise judgments and multi-turn dialogues to reduce self-enhancement bias and align evaluations closer to human standards. Another intriguing avenue is the adoption of Self-Taught Evaluator frameworks, which generate synthetic preference pairs and reasoning traces to iteratively refine model evaluation capabilities Wang et al. (2024f). This approach eliminates dependency on costly human annotations while ensuring that evaluation criteria adapt to evolving tasks and models. The integration of Self-Rewarding Language Models (SRLM) offers another promising path Yuan et al. (2024c). By employing iterative mechanisms like Direct Preference Optimization (DPO), these models continuously improve their instruction-following and reward-modeling capabilities, mitigating issues like reward hacking and overfitting. Building on SRLM, the use of Meta-Rewarding mechanisms introduces a meta-judge role to evaluate and refine judgment quality Wu et al. (2024a). This iterative process addresses biases such as verbosity and positional bias, enhancing both alignment and the ability to evaluate complex tasks. Finally, leveraging synthetic data creation for generating contrasting responses offers a scalable solution for training evaluators Wang et al. (2024f). By iteratively refining evaluations on synthetic preference pairs, models can progressively improve their robustness and adaptability. Combining these approaches with diverse benchmarks Srivastava et al. (2022); Liang et al. (2022), multifaceted evaluation criteria Sellam et al. (2020), and human feedback Xu et al. (2023); Ouyang et al. (2022) can ensure evaluations that are fair, reliable, and aligned with human expectations across various domains.

### v5

@@ 5.3.2 Retrieval-Augmented Generation (RAG)
Recent developments in Retrieval-Augmented Generation (RAG) have explored the capacity of LLMs to self-evaluate and self-improve without the need for annotated datasets or parameter adjustments Chen et al. (2024e). Li and Qiu (2023) introduce the Memory-of-Thought (MoT) framework, a two-stage self-reflective model that enhances an LLM’s reasoning abilities autonomously. In the first stage, the model generates high-confidence reasoning on an unlabeled dataset, storing this as memory. During the testing phase, the model recalls this memory by judging the relevance of each of them with the current question and choose the most relevant ones as demonstration.
@@ 5.4.1 Reasoning Path Selection
Additionally, there are also many works that focus on step-level reasoning path selection, leveraging the judge LLM as the process reward model (PRM) to evaluate state scores. Creswell et al. (2023) decompose the reasoning process into Selection and Inference. In the selection step, they leverage the LLM itself to judge and evaluate each potential reasoning trace, selecting the appropriate one for the following inference step. Xie et al. (2024b) propose the Kwai-STaR framework, which transforms LLMs into state-transition reasoners to judge and select the best reasoning state for themselves in mathematical reasoning. Lightman et al. (2023) train a LLM as the PRM to conduct inference-time supervision and perform best-of-N sampling strategy during reasoning stage. Following them, Setlur et al. (2024) further propose process advantage verifiers (PAVs), generating rewards based on the changes in the likelihood of producing correct responses in the future. Other works simulate advanced cognitive structures as the reasoning process. Hao et al. (2023) employ LLMs as a world model to simulate the state of the environment and perform Monte Carlo Tree Search (MCTS) to improve performance on tasks requiring deliberate path selection. Furthermore, Besta et al. (2024) takes the outputs generated by LLMs as an arbitrary graph. An LLM thought is modeled as a vertex, while an edge is a dependency between thoughts. This framework enables a systematic judge of coherence and logical reasoning for each reasoning state. Yao et al. (2023a) propose Tree-of-Thoughts (ToT), where each thought serves as an intermediate step toward problem-solving. It decomposes reasoning into steps, self-evaluates and judges progress at each state, and uses search algorithms with LMs to judge thought paths through lookahead and backtracking. Besides, there are also studies train critique-based LLM judges Ankner et al. (2024); Yu et al. (2024b); Wang et al. (2024d); Lan et al. ; Xie et al. (2024b); Yao et al. (2024); Yuksel and Sawaf (2024) which provide fine-grained verbal feedback to boost the reasoning process.
@@
DHP Wang et al. (2024i) Various Tasks 400 Human Discernment Score General Performance
EvalBiasBench Park et al. (2024a) Alignment 80 Human Accuracy Various Bias
@@
R-Judge Yuan et al. (2024d) Multi-turn Interaction 569 Human F1, Recall, Spec, Effect Safety
@@ 6.1 General Performance
  • JUDGE-BENCH Tan et al. (2024a), DHP Wang et al. (2024i), RewardBench Lambert et al. (2024), SOS-BENCH Penfever et al. (2024) and JuStRank Gera et al. (2024), which operate at larger scales, utilizing metrics like Cohen’s kappa, Discernment Score, and normalized accuracy to benchmark general LLM performance.
@@ 6.3 Challenging Task Performance
Benchmarks designed for difficult tasks push the boundaries of LLM evaluation. For example, Arena-Hard Auto Li et al. (2024k), JudgeBench Tan et al. (2024a) and Yuan et al. (2024b) select harder questions based on LLMs’ performance for conversational QA and various reasoning tasks, respectively. CALM Ye et al. (2024a) explores alignment and challenging scenarios, using metrics such as separability, agreement, and hacked accuracy to evaluate performance in manually identified hard data sets.
@@ 6.4 Domain-Specific Performance
Domain-specific benchmarks provide task-focused evaluations to assess LLMs’ effectiveness in specialized contexts. Concretely, Raju et al. (2024) measure separability and agreement across tasks, leveraging metrics such as Brier scores for insights in specific domains such as coding, medical, finance, law and mathematics. CodeJudge-Eval Zhao et al. (2024a) specifically evaluates LLMs for judging code generation with execution-focused metrics such as accuracy and F1 score. This idea has also been adopted by several following works in code summarization and generation evaluation Wu et al. (2024b); Yang et al. (2024); Tong and Zhang (2024); Lin et al. (2024); Gomes et al. (2024); Yu et al. (2024e); Sherifi et al. (2024)
@@ 7.1 Bias & Vulnerability
In addition to these general biases, specific evaluative biases emerge when LLMs act as judges. Order Bias is a prominent issue where the sequence of candidates influences preferences Zheng et al. (2023); Wang et al. (2023c); Koo et al. (2023a); Shi et al. (2024b). This bias can distort evaluation outcomes, particularly in pairwise comparisons, and is more pronounced when the quality gap between competing responses is small Shi et al. (2024b); Wang et al. (2023c). Egocentric Bias arises when LLMs favor outputs generated by the same model, compromising objectivity Liu et al. (2023c); Koo et al. (2023a); Wataoka et al. (2024); Panickssery et al. (2024). This issue is particularly pronounced when evaluation metrics are designed using the same model, leading to inflated scores for outputs from the originating model Liu et al. (2023c). Length Bias, another prevalent challenge, skews evaluations by disproportionately favoring longer or shorter responses regardless of quality Zheng et al. (2023); Koo et al. (2023a). Additional biases, such as Misinformation Oversight Bias, Authority Bias, and Beauty Bias, further complicate LLM evaluations. For instance, misinformation oversight bias reflects a tendency to overlook factual errors, authority bias favors statements from perceived authoritative sources, and beauty bias prioritizes visually appealing content over substantive quality Chen et al. (2024b, e); Stephan et al. (2024). Similarly, Verbosity Bias shows a preference for lengthier explanations, often equating verbosity with quality, which can mislead the judgment process Yuan et al. (2024c). Furthermore, Sentiment Bias skews evaluations based on emotional tone, favoring responses with positive phrasing Ye et al. (2024a).
@@
Future Direction. One promising direction for future research is to integrate Retrieval-Augmented Generation (RAG) frameworks into LLM evaluation processes Chen et al. (2024e). By combining generative and retrieval capabilities, these frameworks can reduce biases such as self-preference and factuality issues by grounding evaluations in external, verifiable data sources. Another promising avenue is the use of bias-aware datasets, like OFFSETBIAS, to systematically address inherent biases in LLM-as-a-judge systems Park et al. (2024a). Incorporating such datasets into training pipelines allows LLMs to better distinguish superficial qualities from substantive correctness, thereby enhancing fairness and reliability. Exploring fine-tuned LLMs as scalable judges, exemplified by the JudgeLM framework, represents another intriguing direction Zhu et al. (2023). Techniques like swap augmentation and reference support embedded in this framework can systematically mitigate biases, improve evaluation consistency, and extend the applicability of LLM-based judgments to open-ended tasks. Additionally, advancing zero-shot comparative assessment frameworks offers significant promise Liusie et al. (2023). These frameworks can refine pairwise comparison techniques and implement debiasing strategies, improving fairness and reliability across diverse evaluation domains without the need for extensive prompt engineering or fine-tuning. Finally, methods like JudgeDeceiver-resistant calibration and adversarial phrase detection strategies need further exploration to secure LLM-as-a-judge frameworks from attacks Shi et al. (2024a); Raina et al. (2024); Zhou et al. (2024a); Li et al. (2024d).
@@ 7.2 Dynamic & Complex Judgment
Earlier works about LLM-as-a-judge usually adopt static and straightforward approaches that directly prompt the judge LLM to perform the assessment Zheng et al. (2023). Recently, more dynamic and complex judgment pipelines have been proposed to address various limitations, improving the robustness and effectiveness of LLM-as-a-judge. One approach in direction is to follow the concept of “LLM-as-a-examiner”, where the system dynamically and interactively generates both questions and judgments based on the candidate LLMs’ performance Yu et al. (2024d); Bai et al. (2023a); Kim et al. (2024a). Other works focus on making judgments based on two or more candidate LLMs battle and debating results Moniri et al. (2024); Zhao et al. (2024c); Kenton et al. . These dynamic judgment methods largely improve the judge LLM’s understanding of each candidate and potentially prevent the data contamination problem Deng et al. (2024); Golchin and Surdeanu in LLMs’ evaluation. Additionally, building complex and complicated judgment pipelines or agents is another popular research area Li et al. (2023a); Chan et al. (2023); Zhuge et al. (2024); Zheng et al. (2024). These methods typically involve multi-agent collaboration, along with well-designed planning and memory systems, enabling the judge LLMs to handle more complex and diverse judgment scenarios.
Future Direction. One promising direction for future research is to equip LLMs with human-like judgment capabilities Yuan et al. (2024c); Liang et al. (2024b); Ameli et al. (2024). These designs can borrow insights from human behavior when making a judgment, such as anchoring and comparing, hindsight and reflection, and meta-judgment. Another intriguing avenue would be to develop an adaptive difficulty assessment system using LLMs Hu (2024). This system would adjust the difficulty of questions based on the current performance of the candidates. Such an adaptive and dynamic system could address a significant limitation in LLM evaluation, as static benchmarks often fail to accurately assess LLMs with varying capabilities.
@@ 7.3 Self-Judging
LLM-based evaluators, such as GPT-4, are widely used for assessing outputs but face significant challenges, particularly Egocentric Bias, where models favor their own responses over those from external systems Liu et al. (2023b); Bai et al. (2023a); Zheng et al. (2023). This self-preference undermines impartiality and creates a “chicken or the egg” dilemma: robust evaluators are essential for developing powerful LLMs, yet advancing LLMs depends on unbiased evaluators. Other issues include Self-Enhancement Bias, where models overrate their own outputs Li et al. (2023a), and Reward Hacking, where over-optimization of specific signals leads to less generalizable evaluations Wu et al. (2024a). Additionally, reliance on Static Reward Models limits adaptability, while biases like Positional and Verbosity Biases distort judgments by favoring response order or length over quality Yuan et al. (2024e); Wang et al. (2024h). The high cost and limited scalability of human annotations further complicate the creation of dynamic and reliable evaluation systems Srivastava et al. (2022); Liang et al. (2022).
@@
Future Direction. One promising direction for future research is the development of collaborative evaluation frameworks like Peer Rank and Discussion (PRD) Li et al. (2023a). These frameworks leverage multiple LLMs to collectively evaluate outputs, using weighted pairwise judgments and multi-turn dialogues to reduce self-enhancement bias and align evaluations closer to human standards. Another intriguing avenue is the adoption of Self-Taught Evaluator frameworks, which generate synthetic preference pairs and reasoning traces to iteratively refine model evaluation capabilities Wang et al. (2024h). This approach eliminates dependency on costly human annotations while ensuring that evaluation criteria adapt to evolving tasks and models. The integration of Self-Rewarding Language Models (SRLM) offers another promising path Yuan et al. (2024e). By employing iterative mechanisms like Direct Preference Optimization (DPO), these models continuously improve their instruction-following and reward-modeling capabilities, mitigating issues like reward hacking and overfitting. Building on SRLM, the use of Meta-Rewarding mechanisms introduces a meta-judge role to evaluate and refine judgment quality Wu et al. (2024a). This iterative process addresses biases such as verbosity and positional bias, enhancing both alignment and the ability to evaluate complex tasks. Finally, leveraging synthetic data creation for generating contrasting responses offers a scalable solution for training evaluators Wang et al. (2024h). By iteratively refining evaluations on synthetic preference pairs, models can progressively improve their robustness and adaptability. Combining these approaches with diverse benchmarks Srivastava et al. (2022); Liang et al. (2022), multifaceted evaluation criteria Sellam et al. (2020), and human feedback Xu et al. (2023); Ouyang et al. (2022) can ensure evaluations that are fair, reliable, and aligned with human expectations across various domains.